LINE INTENSITIES OF CH_{3}OH^{a} AT 400 cm^{−}^{1}

\underline{CAROLYN S. BRAUER}, KEEYOON SUNG, JOHN C. PEARSON, LINDA R. BROWN, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, CA 91109; LI-HONG XU, Centre for Laser, Atomic and Molecular Sciences (CLAMS), Department of Physical Sciences, University of New Brunswick, Saint John, N.B., Canada E2L 4L5.

Line intensities of methanol (CH_{3}OH) have been measured between 300 and 500 cm^{−1} . This room temperature absorption spectrum was recorded with a Bruker IFS 125 HR Fourier transform spectrometer located at the Jet Propulsion Laboratory, at resolutions between 0.0014 and 0.0050 cm^{−1}, using a 20.1 cm glass cell with wedged high density polyethylene windows. This work has been undertaken to investigate the apparent incongruity in rotational temperature diagrams derived from astronomical observations of methanol, which may be due to not including the vibration-, torsion- and K-dependence of the dipole moment in the rotational intensities. The results from this study will support the analyses of astronomical observations taken from orbit by the Herschel HIFI instrument.

aResearch described in this paper was performed at the Jet Propulsion Laboratory, California Institute of Technology, under contracts and cooperative agreements with the National Aeronautics and Space Administration.